# SEAFormer: A Spatial Proximity and Edge-Aware Transformer for Real-World Vehicle Routing Problems

### v1

@@ 1 Introduction
Recent efforts to address the scalability limitations of autoregressive models introduce new challenges when applied to RWVRPs. Divide-and-conquer methods (Hou et al., 2022; Nasehi et al., 2025; Zheng et al., 2024) cluster customers first, then solve each cluster independently. However, in RWVRPs this creates a fundamental circular dependency: assessing whether a cluster is feasible requires knowing the vehicle’s state upon arrival (e.g., remaining battery in EVRPCS or current time in VRPTW), but that state depends on the full route, which is not yet known during clustering. This mismatch often leads to infeasible clusters that require costly post-hoc repair or prevents the discovery of high-quality routes altogether (refer to Appendix G). Other scalable architectures (Gao et al., 2024; Luo et al., 2023, 2025a) face their own limitations, including substantial computational overhead or reliance on supervised learning which requires a set of pre-computed optimal solutions that are themselves computationally expensive to obtain.
@@ 2 Related work
We review existing learning-based approaches to RWVRPs and VRP, focusing on why they struggle to simultaneously achieve solution quality, computational efficiency, and constraint satisfaction at real-world scales. A more comprehensive discussion is provided in Appendix K.
@@ 2.2 Scalable neural approaches
Recent approaches for scalable VRP solutions (Hou et al., 2022; Zheng et al., 2024; Nasehi et al., 2025; Ye et al., 2024) are based on divide-and-conquer method. Such techniques, however, cannot solve RWVRPs effectively as they cluster nodes before determining visit sequences, yet feasibility of a solution for RWVRPs depends on those sequences. Whether a vehicle can serve a cluster requires knowing its state upon arrival, such as battery level in EVRPCS or elapsed time in VRPTW, which is unavailable during clustering. This circular dependency causes clusters to violate constraints once sequenced, requiring expensive repairs or prevents finding high-quality solutions.
@@ 3 Clustered Proximity Attention
This rotation softens the deterministic cluster boundaries that would otherwise push nearby nodes apart, keeping close nodes together and improving transitions between clusters. After computing the partitioning scores $S=\{S_{\alpha_{1}},S_{\alpha_{1}}^{\prime},S_{\alpha_{2}},$...\} across different proximity weights and rotations, we partition nodes into clusters of size $M$ using the partitioning score. Note that when $n$ is not perfectly divisible by (i.e., when $\lceil n/M\rceil\times M>n$, we pad the final cluster with depot nodes to ensure uniform cluster size $M$ across all partitions).
@@ 4 SEAFormer Architecture
Figure 2: SEAFormer Architecture. The dual-module encoder embeds nodes through CPA and edges through a lightweight residual module. The dual-path decoder combines edge-aware guidance (heatmap) with sequential node selection (attention), unified through logit fusion. This design enables scalable training on large number of nodes while handling diverse RWVRP constraints.
@@
For problem variants with optional service nodes such as EVRPCS (charging stations) and VRPRS (replenishment stops), we augment the spatial encoder with a specialized optional node processing pathway. This parallel encoding layer generates embeddings for optional nodes $H_{o}\in\mathbb{R}^{f\times d}$, where $f$ represents the number of service facilities. These embeddings are then fused with customer embeddings through a learnable mechanism, and the combined representations pass through the same batch normalization and feed-forward layers. This allows each customer embedding to account for nearby optional facilities and their impact on route feasibility, which is crucial in problems where strategic service stops can substantially enhance or even enable high-quality solutions. A detailed description of optional node embedding is given in Appendix H.
@@ 5 Experiments
Benchmarks. We evaluate SEAFormer across four datasets: (i) Random RWVRP benchmarks, comprising 100 instances per problem size with up to 7K nodes (results up to 1k are shown in table 1, and larger values in appendix C) following (Kwon et al., 2020; Zhou et al., 2023); (ii) Random VRP benchmarks with up to 7K nodes following (Zhou et al., 2023); (iii) real-world VRP, EVRPCS, and VRPTW datasets; (iv) cross-distribution datasets (Zhou et al., 2023).
Implementation. For all problems, we train the model for 2000 epochs on 100-customer instances, then 200 epochs on 500-customer instances, and 100 epochs on 1000-customer instances. As in prior work, problem instances are uniformly sampled from the $0,1]^{2}$ space, and demands drawn from a discrete uniform distribution on $[1,10]$. Asymmetric VRPs are created using our dataset generation procedure (refer to appendix A.4.1). The Adam optimizer (Kingma, 2014) is employed for training, with detailed hyperparameters for each problem instance as well as training times provided in Appendix L and Appendix M.
@@
Baselines. We compare SEAFormer with 1) Classical Solvers: HGS-PyVRP (Wouda et al., 2024), LKH (Helsgaun, 2017), and OR-Tools; 2) Construction-based NCO Methods: POMO (Kwon et al., 2020), MTPOMO (Liu et al., 2024), EVRPRL (Lin et al., 2021), EVGAT (Wang et al., 2024), LEHD (Luo et al., 2023), RELD (Huang et al., 2025), ELG (Gao et al., 2024) L2C-insert (Luo et al., 2025b), UniteFormer (Meng et al., ), Eformer (Meng et al., 2025) DAR (Wang et al., 2025), BLEHD (Luo et al., 2025a), and RouteFinder (Berto et al., 2025); 3) Divide-and-conquer based methods: GLOP (Ye et al., 2024), DeepMDV (Nasehi et al., 2025), and UDC (Zheng et al., 2024). Not all existing methods apply to every RWVRP variant, so comparisons are limited to applicable approaches. For details on baselines and their integration, see Appendix N.
Table 1: Objective function (Obj.), Gap to the OR-tools (Gap), and solving time (Time) on 100, 500, and 1,000-node RWVRPs. All test sets contain 100 instances following settings in (Zhou et al., 2023). The overall best performance is in bold and the best learning method is marked by shade. OR-Tools results are not optimal, as execution was stopped early due to time limits. Methods not tailored to EVRPCS or VRPRS are excluded due to sub-optimal performance. (Refer to Section G).
@@ Appendix I Visual Example of CPA
Appendix K Additional related work
K.1 Classical approaches and their limitations
@@
K.2 Learning-based approaches
The intersection of machine learning and combinatorial optimization has yielded approaches that directly produce high-quality VRP solutions without iterative refinement. These methods fall into two architectural paradigms: (1) Autoregressive models that build solutions sequentially, adding one decision at a time (Khalil et al., 2017; Kool et al., 2018; Kwon et al., 2020; Hou et al., 2022; Luo et al., 2025a, 2023; Zheng et al., 2024; Berto et al., 2025; Nasehi et al., 2025), and (2) Non-autoregressive models that generate complete solutions simultaneously, typically through learned heatmaps (Nowak et al., 2018; Kool et al., 2022; Ye et al., 2024; Xiao et al., 2024). Training paradigms include supervised learning from optimal solutions or reinforcement learning to directly optimize solution quality (Joshi et al., 2019).
@@
Divide-and-conquer strategies have emerged as the dominant approach for large-scale instances (Nowak et al., 2018; Li et al., 2021; Zong et al., 2022; Fu et al., 2021; Zheng et al., 2024; Nasehi et al., 2025). TAM (Hou et al., 2022) employs two-stage decomposition followed by LKH3 (Helsgaun, 2017) for sub-problem resolution. GLOP (Ye et al., 2024) integrates global partitioning through non-autoregressive models with local autoregressive construction. UDC (Zheng et al., 2024) addresses sub-optimal partitioning through robust training procedures, combining GNN-based decomposition with specialized sub-problem solvers.
@@
K.3 Local attention mechanism
@@
Appendix M Model parameters and training time
@@
Appendix N Baselines, codes and licenses
@@
DeepMDV (Nasehi et al., 2025) https://github.com/SaeedNB/DeepMDV MIT License
@@
Appendix O MDVRP extensibility
@@
Appendix P Use of Large Language Models

### v2

@@ 1 Introduction
Recent efforts to address the scalability limitations of autoregressive models introduce new challenges when applied to RWVRPs. Divide-and-conquer methods (Hou et al., 2022; Nasehibasharzad et al., 2025; Zheng et al., 2024) cluster customers first, then solve each cluster independently. However, in RWVRPs this creates a fundamental circular dependency: assessing whether a cluster is feasible requires knowing the vehicle’s state upon arrival (e.g., remaining battery in EVRPCS or current time in VRPTW), but that state depends on the full route, which is not yet known during clustering. This mismatch often leads to infeasible clusters that require costly post-hoc repair or prevents the discovery of high-quality routes altogether (refer to Appendix G). Other scalable architectures (Gao et al., 2024; Luo et al., 2023, 2025a) face their own limitations, including substantial computational overhead or reliance on supervised learning which requires a set of pre-computed optimal solutions that are themselves computationally expensive to obtain.
@@ 2 Related work
We review existing learning-based approaches to RWVRPs and VRP, focusing on why they struggle to simultaneously achieve solution quality, computational efficiency, and constraint satisfaction at real-world scales. A more comprehensive discussion is provided in Appendix J.
@@ 2.2 Scalable neural approaches
Recent approaches for scalable VRP solutions (Hou et al., 2022; Zheng et al., 2024; Nasehibasharzad et al., 2025; Ye et al., 2024) are based on divide-and-conquer method. Such techniques, however, cannot solve RWVRPs effectively as they cluster nodes before determining visit sequences, yet feasibility of a solution for RWVRPs depends on those sequences. Whether a vehicle can serve a cluster requires knowing its state upon arrival, such as battery level in EVRPCS or elapsed time in VRPTW, which is unavailable during clustering. This circular dependency causes clusters to violate constraints once sequenced, requiring expensive repairs or prevents finding high-quality solutions.
@@ 3 Clustered Proximity Attention
This rotation softens the deterministic cluster boundaries that would otherwise push nearby nodes apart, keeping close nodes together and improving transitions between clusters. After computing the partitioning scores $S=\{S_{\alpha_{1}},S_{\alpha_{1}}^{\prime},S_{\alpha_{2}},$...\} across different proximity weights and rotations, we partition nodes into clusters of size $M$ using the partitioning score. Note that when $n$ is not perfectly divisible by (i.e., when $\lceil n/M\rceil\times M>n$), the last cluster is padded with depot nodes to maintain size $M$ for all clusters.
@@ 4 SEAFormer Architecture
Figure 2: SEAFormer Architecture. The dual-module encoder embeds nodes via CPA and edges via a lightweight residual module. The dual-path decoder combines edge-aware guidance (heatmap) with sequential node selection (attention) through logit fusion. This design enables scalable training on large node sets and diverse RWVRP constraints. A visual illustration of CPA is provided in Appendix I.
@@
For problem variants with optional service nodes such as EVRPCS (charging stations) and VRPRS (replenishment stops), we augment the spatial encoder with a specialized optional node processing pathway. This parallel encoding layer generates embeddings for optional nodes $H_{o}\in\mathbb{R}^{f\times d}$, where $f$ represents the number of service facilities. These embeddings are then fused with customer embeddings through a learnable mechanism, and the combined representations pass through the same batch normalization and feed-forward layers. This allows each customer embedding to account for nearby optional facilities and their impact on route feasibility, which is crucial in problems where strategic service stops can substantially enhance or even enable high-quality solutions. A detailed description of optional node embedding is presented in Appendix H.
@@ 5 Experiments
Benchmarks. We evaluate SEAFormer across four datasets: (i) Random RWVRP benchmarks, comprising 100 instances per problem size with up to 7K nodes (results up to 1k are shown in table 1, and larger values in appendix C) following (Kwon et al., 2020; Zhou et al., 2023); (ii) Random VRP benchmarks with up to 7K nodes following (Zhou et al., 2023); (iii) real-world VRP, EVRPCS (Mavrovouniotis et al., 2020), and VRPTW (Gehring and Homberger, ) datasets; (iv) cross-distribution datasets (Zhou et al., 2023).
Implementation. For all problems, we train the model for 2000 epochs on 100-customer instances, then 200 epochs on 500-customer instances, and 100 epochs on 1000-customer instances. As in prior work, problem instances are uniformly sampled from the $0,1]^{2}$ space, and demands drawn from a discrete uniform distribution on $[1,10]$. Asymmetric VRPs are created using our dataset generation procedure (refer to appendix A.4.1). The Adam optimizer (Kingma, 2014) is employed for training, with detailed hyperparameters for each problem instance as well as training times provided in Appendix K and Appendix L.
@@
Baselines. We compare SEAFormer with 1) Classical Solvers: HGS-PyVRP (Wouda et al., 2024), LKH (Helsgaun, 2017), and OR-Tools; 2) Construction-based NCO Methods: POMO (Kwon et al., 2020), MTPOMO (Liu et al., 2024), EVRPRL (Lin et al., 2021), EVGAT (Wang et al., 2024), LEHD (Luo et al., 2023), RELD (Huang et al., 2025), ELG (Gao et al., 2024) L2C-insert (Luo et al., 2025b), UniteFormer (Meng et al., ), Eformer (Meng et al., 2025) DAR (Wang et al., 2025), BLEHD (Luo et al., 2025a), and RouteFinder (Berto et al., 2025); 3) Divide-and-conquer based methods: GLOP (Ye et al., 2024), DeepMDV (Nasehibasharzad et al., 2025), and UDC (Zheng et al., 2024). Not all existing methods apply to every RWVRP variant, so comparisons are limited to applicable approaches. For details on baselines and their integration, see Appendix M.
Table 1: Objective function (Obj.), Gap to the OR-tools (Gap), and solving time (Time) on 100, 500, and 1,000-node RWVRPs. All test sets contain 100 instances following settings in (Zhou et al., 2023). The overall best performance is in bold and the best learning method is marked by shade. OR-Tools results are not optimal, as execution was stopped early due to time limits. Methods not tailored to EVRPCS or VRPRS are excluded due to sub-optimal performance (Refer to Section G).
@@ Appendix I Visual Example of CPA
Appendix J Additional related work
J.1 Classical approaches and their limitations
@@
J.2 Learning-based approaches
The intersection of machine learning and combinatorial optimization has yielded approaches that directly produce high-quality VRP solutions without iterative refinement. These methods fall into two architectural paradigms: (1) Autoregressive models that build solutions sequentially, adding one decision at a time (Khalil et al., 2017; Kool et al., 2018; Kwon et al., 2020; Hou et al., 2022; Luo et al., 2025a, 2023; Zheng et al., 2024; Berto et al., 2025; Nasehibasharzad et al., 2025), and (2) Non-autoregressive models that generate complete solutions simultaneously, typically through learned heatmaps (Nowak et al., 2018; Kool et al., 2022; Ye et al., 2024; Xiao et al., 2024). Training paradigms include supervised learning from optimal solutions or reinforcement learning to directly optimize solution quality (Joshi et al., 2019).
@@
Divide-and-conquer strategies have emerged as the dominant approach for large-scale instances (Nowak et al., 2018; Li et al., 2021; Zong et al., 2022; Fu et al., 2021; Zheng et al., 2024; Nasehibasharzad et al., 2025). TAM (Hou et al., 2022) employs two-stage decomposition followed by LKH3 (Helsgaun, 2017) for sub-problem resolution. GLOP (Ye et al., 2024) integrates global partitioning through non-autoregressive models with local autoregressive construction. UDC (Zheng et al., 2024) addresses sub-optimal partitioning through robust training procedures, combining GNN-based decomposition with specialized sub-problem solvers.
@@
J.3 Local attention mechanism
@@
Appendix L Model parameters and training time
@@
Appendix M Baselines, codes and licenses
@@
DeepMDV (Nasehibasharzad et al., 2025) https://github.com/SaeedNB/DeepMDV MIT License
@@
Appendix N MDVRP extensibility
@@
Appendix O Use of Large Language Models